Diffusion Morning Report 11/22/02

Ted Berg tberg at snf.stanford.edu
Fri Nov 22 07:32:13 PST 2002


Good Morning all,

Tylan 2s TCA controller seems to be working OK, Dichlorosilane seems to
be holding after change, LTO will be going down next week for clean.


Date: 11/21/2002
Tube Name/Number Status Comments THICKNESS-Change at
Tylan 1 Up
Tylan 2 Up
Tylan 3 Up
Tylan 4 Up
Tylan 5-BBr3 Up
Tylan 6-POCl3 Up
Tylan 7-Dirty Ox Up
Tylan 8-SiGe up  6.5?
Tylan 9-Poly Up  18.3/30
Tylan 10-Nitride Up   11.9/30
Tylan 11-BPSG-LTO yellow Uniformity issues 24.2/30
Tylan 12-FGA Up
Tystar Up  13.1/30???
Doped Poly Down
TEOS yellow  testing again
Tylan 16 N/A
Burn Box Pressure OK Checking this Morning

GAAS 22 UP
GAAS 23 UP

SiH4 for furnaces   Change at 100PSI  Currently @490 PSI

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20021122/97b64feb/attachment.html>


More information about the diffusion mailing list